### v1

# ComputerRL: Scaling End-to-End Online Reinforcement Learning for Computer Use Agents

Abstract
We introduce ComputerRL, a framework for autonomous desktop intelligence that enables agents to operate complex digital workspaces skillfully. ComputerRL features the API-GUI paradigm, which unifies programmatic API calls and direct GUI interaction to address the inherent mismatch between machine agents and human-centric desktop environments. Scaling end-to-end RL training is crucial for improvement and generalization across diverse desktop tasks, yet remains challenging due to environmental inefficiency and instability in extended training. To support scalable and robust training, we develop a distributed RL infrastructure capable of orchestrating thousands of parallel virtual desktop environments to accelerate large-scale online RL. Furthermore, we propose Entropulse, a training strategy that alternates reinforcement learning with supervised fine-tuning, effectively mitigating entropy collapse during extended training runs. We employ ComputerRL on open models GLM-4-9B-0414 and Qwen2.5-14B, and evaluate them on the OSWorld benchmark. The AutoGLM-OS-9B based on GLM-4-9B-0414 achieves a new state-of-the-art accuracy of 48.1%, demonstrating significant improvements for general agents in desktop automation. The algorithm and framework are adopted in building AutoGLM (Liu et al., 2024a).
(a) The success rates of agents on OSWorld.
(b) ComputerRL training reward curves (95% CIs).
Figure 1: ComputerRL enables efficient end-to-end online policy optimization for OS agents. (a) On OSWorld (Xie et al., 2024), AutoGLM-OS, trained with ComputerRL, outperforms state-of-the-art agents. (b) Our Entropulse approach yields higher average training rewards and improves both learning efficiency and final performance over conventional methods.
Figure 2: Examples of AutoGLM-OS’s execution on four user tasks, including image processing between GIMP and LibreOffice Writer, monitoring system resource usage in Terminal, table calculation in LibreOffice Calc, and document formatting in LibreOffice Writer.
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; GLM et al., 2024; Zeng et al., 2022; Team et al., 2023; DeepSeek-AI et al., 2025; Bai et al., 2023a; Yang et al., 2025) have dramatically expanded the scope and depth of artificial intelligence (AI) capabilities, driving a profound re-examination of our understanding of machine intelligence. Among all scenarios, the emergence of LLM-based GUI (graphical user interfaces) agents, capable of independently perceiving, reasoning, and executing complex tasks on user devices, has aroused particular interests from researchers (Xi et al., 2023; Wang et al., 2023; Liu et al., 2023). Given that desktops remain central to intelligence-intensive tasks, developing effective computer use agents is crucial for fundamentally transforming human-computer interactions (Agashe et al., 2025; Wu et al., 2024) and lifting capabilities of AI to higher levels.
Despite previous attempts to develop computer use agents (Agashe et al., 2025; Lei et al., 2024; Xie et al., 2025), enabling them to operate autonomously over extended periods in real-world scenarios remains a significant challenge. The first primary obstacle arises from the fact that GUIs are inherently designed for human interaction, making the simulation of human actions by GUI agents (Liu et al., 2024b; OpenAI, 2025; Qin et al., 2025) a non-trivial and cumbersome endeavor. Second, current mainstream approaches of behavior cloning (BC) (Bain & Sammut, 1995; Bratko et al., 1995), including manual annotation (He et al., 2025) and model distillation (Sun et al., 2024; Xu et al., 2024), are limited in scalability and effectiveness. Manual annotation, while precise, is prohibitively labor-intensive for complex tasks. Model distillation, on the other hand, is constrained by the performance of the teacher models, limiting overall capability. Both methods typically exhibit poor generalization and limited error recovery abilities. Finally, although reinforcement learning (RL) has shown potential for desktop automation tasks (Lu et al., 2025; Feng et al., 2025), its practical application remains restricted due to computational complexity and methodological challenges. Complex environment, slow convergence, and known inefficiencies in RL training (Xie et al., 2024; Bonatti et al., 2024; Yu et al., 2025; Fu et al., 2025) severely limit its large-scale adoption in training computer use agents.
In this work, we propose ComputerRL, an end-to-end algorithmic framework designed to advance desktop-level planning, reasoning, and device operation, including a new API-GUI interaction paradigm, a scalable RL training infrastructure for computer environment, and a RL algorithm for extended effective training. First, we introduce API-GUI, a large-scale, automatically constructed API ecosystem that enables the agent to transcend the inherent biases of human-oriented operational paradigms. It instead leverages a more machine-oriented approach for device interaction, which combines API calls and GUI actions and thereby significantly enhances both the versatility and overall performance of the agent. Second, we develop a distributed training infrastructure utilizing virtual machine clusters based on Docker and gRPC protocols for scalability, which is fully compatible with AgentBench (Liu et al., 2023). This infrastructure supports thousands of parallel environments, ensuring high scalability and consistent interactions across all environments. Additionally, we combine the training infrastructure with the AgentRL framework (Zhang et al., 2025) for efficient asynchronous training, substantially accelerating the training process. Finally, to counteract stagnation and convergence issues in extended RL training—specifically, entropy collapse and rising KL divergence—we propose Entropulse, which alternates between RL and periodic SFT phases. This approach maintains exploratory capacity and ensures continuous performance gains (Figure 1(b)).
As a result, by harnessing end-to-end RL and optimization in the desktop environment, ComputerRL has achieved remarkable improvement in understanding and operating GUIs. Evaluation on the OSWorld benchmark (Xie et al., 2024) shows ComputerRL’s significant improvements (see Figure 1(a)) in computer use challenges, achieving a success rate of 48.1% (with 64% performance gain from RL), outperforming other state-of-the-art models including OpenAI CUA o3 (42.9%), UI-TARS-1.5 (42.5%), and Anthropic Claude Sonnet 4 (30.7%).
In summary, our contributions are as follows:
  • We propose a new interaction paradigm, shift from human-centric to machine-oriented interaction frameworks by introducing a large-scale, automatically constructed API ecosystem integrated with conventional GUI operations. This approach addresses the inherent mismatch between human-designed interfaces and artificial agent capabilities, while achieving superior operational efficiency and generalization performance on computer-based tasks.
  • We establish a large-scale, distributed RL infrastructure for computer use agents by reconstructing virtual machine clusters, achieving unprecedented scalability with thousands of parallel environments and seamless AgentBench compatibility, thereby overcoming the critical bottleneck that has limited RL-based computer use agent training to scale experiments and enabling breakthrough results in large-scale agent training.
  • We introduce Entropulse, a novel training methodology that systematically addresses the challenges of entropy collapse and KL divergence accumulation in extended RL training through strategic alternation between RL and SFT phases, enabling sustained performance improvements and achieving state-of-the-art performance on the OSWorld benchmark.
2 The ComputerRL Framework
The human-oriented design of GUI environments leads to inefficient interactions for agents, which impedes effective operation and learning. Furthermore, limited environment scalability poses a significant barrier to large-scale agent training. In this section, we present the design and implementation of the ComputerRL framework, as in Figure 3. We propose an API-GUI paradigm powered by LLMs, which integrates human-like GUI interactions with efficient API invocation. Furthermore, we develop a robust and highly scalable Ubuntu Desktop environment designed for large-scale parallel experimentation. To facilitate efficient model training, we use a fully asynchronous RL framework (Zhang et al., 2025).
Figure 3: Overview of ComputerRL framework. We introduce an API-GUI action paradigm that seamlessly integrates automatically constructed APIs with GUI actions to improve agent efficiency and effectiveness. A large-scale parallel desktop environment with over 1,000 real-world instances, combined with an asynchronous RL framework, enables efficient sampling and robust agent training.
2.1 General API-GUI Paradigm
A significant challenge confronted by existing GUI agents arises from their reliance on human-like GUI interactions for device manipulation, which poses substantial difficulties for GUI agents. Conversely, employing a diverse set of APIs for application control offers a more efficient and agent-friendly means of operation. However, this approach introduces its own set of challenges, including the complexity of API implementation, limitations in agent adaptability and flexibility, and the fact that many applications restrict CLI access due to security considerations. To address these issues, we propose an innovative API-GUI paradigm that unifies both action spaces. This design enables agents to leverage the efficiency of API invocation while retaining the versatility of GUI-based operations for general tasks.
Moreover, we develop an automated workflow for desktop application API development. Leveraging LLMs, our framework automates key steps, including requirement analysis, API implementation, and generation of basic test cases. The workflow significantly lowers the barrier for users to create APIs tailored to various applications, thereby facilitating rapid development and deployment.
API Construction In this section, we introduce a methodology for leveraging LLMs to automate API construction. We propose an automated workflow wherein users need only supply exemplar tasks performed within the target application; our LLM then autonomously generates both the necessary API code and corresponding test cases. The workflow comprises three primary stages: requirement analysis, API implementation, and test case generation.
  • Requirement Analysis During the requirements analysis phase, users provide a set of task examples related to the target application as input. The workflow leverages our LLM to analyze these task instances, extracting the essential functionalities required for task completion. It then compares these requirements against the existing API interface definitions to identify potential gaps. If uncovered functionalities are detected, the system automatically generates new API interfaces along with their corresponding parameter specifications.
    Notably, we limit the generated interfaces to encapsulate only general-purpose functionalities, thereby avoiding excessive complexity and the proliferation of APIs. This design choice mitigates implementation difficulty and reduces the adaptation burden on the agent.
  • API Implementation Upon obtaining the interface definitions, the workflow systematically iterates over each interface and its associated parameters. For each specification, it leverages the designated Python libraries of the target application to implement the corresponding API functionalities. Additionally, the workflow incorporates error-handling mechanisms and logging to facilitate human debugging and future maintenance. This automated approach not only streamlines API development but also enhances consistency and reusability across different application contexts.
  • Test Case Generation Following the implementation of API functionalities, the workflow conducts fundamental unit testing to ensure the correctness and robustness of each API. Specifically, the testing process verifies: (1) whether the API can be invoked without runtime errors, and (2) whether the API returns correct results across a range of parameter inputs. For API implementations that fail these tests, the workflow provides detailed error feedback to the API implementation module, which then autonomously attempts corrections until the APIs pass all tests.
This automated methodology substantially lowers the manual effort required, enabling the creation of application-specific API sets with minimal human intervention. As a result, the barrier for users to develop APIs for diverse applications is significantly reduced. We have developed API sets for multiple commonly used default applications in Ubuntu and integrated them into our Ubuntu virtual machine environment, where their effectiveness has been empirically validated through experiments.
2.2 Stable Ubuntu Environment for Large-Scale Parallelism
A stable and highly scalable Ubuntu environment is fundamental for constructing behavior cloning datasets and facilitating large-scale RL training. Building on OSWorld (Xie et al., 2024), we have identified several shortcomings of its native implementation. The key challenges identified with the vanilla OSWorld are as follows:
  • Resource Intensiveness and Stability: Virtual machines launched via OSWorld are highly CPU-intensive and consume substantial compute resources. High concurrency on a single node often leads to significant instability, including severe performance degradation and frequent system freezes.
  • Network Bottlenecks: Under heavy workloads, the environment exhibits considerable network communication overhead, resulting in frequent connection failures and even loss of IP addresses. Such network issues critically hinder agent-environment interaction and data logging.
  • Lack of Native Distributed Support: OSWorld does not natively support multi-node clustering, prohibiting efficient distributed deployment and operation across compute clusters.
To address these limitations, we build a robust and massively parallelizable OSWorld infrastructure (see Figure 3), with the following design innovations:
  • Standardized, Decoupled Interface: The environment is refactored via the AgentBench API, providing a unified, modular interface. This abstraction layer enables seamless integration with any learning algorithm or agent architecture, and naturally decouples environment execution from the computational back-end. Consequently, resources can be managed more flexibly, and agents can interact with the environment irrespective of the underlying hardware cluster.
  • Lightweight & Efficient VM Deployment: Our solution utilizes qemu-in-docker  [^1] https://github.com/qemus/qemu to orchestrate Ubuntu VM instances in a containerized manner. We build streamlined VM images that mitigate prior network stack issues, bundle essential utilities directly into the image, and reduce resource requirements through optimizing unneeded services or settings. This significantly reduces per-instance CPU usage, enabling a greater density of concurrent environments.
  • Distributed Multi-Node Clustering: To achieve genuine scalability, we employ a gRPC-based  [^2] https://github.com/grpc inter-process communication protocol that links multiple CPU nodes into a distributed cluster. A unified controller service centrally manages resource allocation, environment orchestration, and inter-node synchronization.
  • Web-based Visualization and Monitoring: Our controller incorporates a versatile web interface, enabling real-time visualization of environment statuses, agent states, and cluster resource allocation. The module dramatically improves usability, operational transparency, and facilitates rapid debugging during large-scale experiments.
Through these technical improvements, our system supports deployment of serveral thousands of concurrent environments on a multi-node CPU cluster, as validated by extensive empirical evaluation. Results confirm our platform’s superior stability, resource efficiency, and scalability, making it an enabling infrastructure for large-scale RL and agent-based research.
2.3 Full-Asynchronous RL Framework for Efficient Training
In addition to ensuring a stable and scalable environment, an efficient RL training framework is crucial for supporting large-scale model training. Most existing RL frameworks rely on synchronous training paradigms, where rollout collection and parameter updates proceed in alternating phases, resulting in substantial training inefficiencies. To address these limitations, we utilize AgentRL (Zhang et al., 2025) for fully asynchronous RL training with the following core design principles:
  • Dedicated Resource Partitioning: Data collection runs on its resource pool, separated from training. The trainer streams data from the replay engine, so data ingestion and computation do not block each other.
  • Dynamic Batch Sizing: The trainer processes incoming data as soon as enough is available, using flexible batch sizes. This reduces idle time and improves resource efficiency.
  • Modular Component Isolation: Actor, reference, and critic modules run independently from the replay engine, each with dedicated resources. We utilize PyTorch distributed groups and NCCL for efficient parameter sharing, which stabilizes GPU memory usage and enhances throughput.
  • Off-policy Bias Mitigation: We limit the replay buffer size and sync all collected trajectories to the training group after each update. This ensures that each trajectory remains close to the latest policy, thereby minimizing off-policy bias in practice.
Through a stable, high-concurrency desktop environment and the decoupling of training from rollout, we markedly enhance the efficiency of sampling and RL training. Our system achieves a high average power consumption per GPU, reflecting optimal resource utilization. This design supports scalable, high-throughput RL training by enabling dynamic workload balancing, leading to significant improvement in hardware efficiency and overall training throughput.
3 The ComputerRL Training
Building upon the stable and efficient environment and training framework 2, we establish a robust foundation for large-scale agent training. However, scaling end-to-end training is hindered by challenges in initializing a capable and diverse base policy and by entropy collapse during RL, which limits sustained improvement. In this section, we detail our approach to scalable ComputerRL training, highlighting how these algorithmic innovations facilitate efficient learning in complex desktop environments.
Figure 4: Overview of ComputerRL training. It includes three stages: (1) BC cold-start with trajectories collected via general LLMs; (2) RL with step-level GRPO using verifiable, rule-based rewards; (3) Entropulse, which alternates RL with SFT on successful rollouts to restore entropy and sustain learning.
3.1 Behavior Cloning Setup
To perform a cold start for our model, we employ BC as the initial stage of training. By imitating user interactions, BC enables agents to acquire foundational competencies, thereby facilitating rapid adaptation to computer operations and tasks.
Trajectory Collection with Multiple General LLMs. We manually collect extensive, well-defined tasks and corresponding evaluation functions. However, the large-scale collection of high-quality interaction trajectories remains a fundamental challenge. Manual annotation, while precise, is prohibitively expensive in terms of both human labor and the overhead of developing and maintaining dedicated annotation tools. Furthermore, relying solely on a single model for trajectory generation inevitably results in a limited and homogeneous data distribution, with the attainable quality constrained by the upper bound of that model’s capabilities.
To address these limitations, we leverages the complementary strengths of multiple advanced models to collect a diverse and high-quality set of interaction trajectories. Concretely, our data pipeline consists of three key stages:
  1. 1.
    Initial Sampling: For each task curated as described, we utilize closed-source LLMs to sample several trajectories per task independently. We record both the complete interaction trajectories and the outputs produced by the respective evaluation functions. This procedure yields a rich set of diverse trajectories that serve as the foundation for subsequent data augmentation and model adaptation.
  2. 2.
    Task-Oriented Fine-Tuning with Stratified Sampling: Following initial data collection, we perform a stratified analysis of task outcomes by categorizing all tasks into three groups based on achieved accuracy:
    • Fully Solved ($\mathrm{acc}=100\%$),
    • Partially Solved ($0\%<\mathrm{acc}<100\%$),
    • Unsolved ($\mathrm{acc}=0\%$).
  3. 3.
    Trajectory Augmentation for Hard Tasks: For partially solved tasks, we conduct supervised fine-tuning of our backbone model using the initial trajectories as input. The fine-tuned model is then employed to sample more trajectories for each task, thus substantially expanding the coverage and quality of trajectories for tasks where model proficiency was previously limited.
    For tasks classified as unsolved, we build a model pool consists of high-performing models and randomly pick a model to determine the following action at each step in the interaction process. This approach is motivated by the empirical observation that, while aggregate model performances are comparable, considerable inter-model variance exists at the task level, reflecting distinct areas of expertise. Model collaboration thus enables the generation of trajectories that would be unattainable using any single model in isolation.
We systematically aggregate and filter the collected interaction data, retaining only successful trajectories, and employ them for supervised fine-tuning of the model. This strategy equips the model with robust desktop manipulation capabilities and foundational reasoning abilities, significantly enhancing the performance of the base model.
3.2 Reinforcement Learning from Verifiable Rewards
Step-Level Group Relative Policy Optimization. We have extended the GRPO algorithm (Shao et al., 2024) to the step-level, making it more suitable for agent RL training. For each task $\tau$, the policy $\pi_{\theta}$ interacts with the desktop environment and samples $G$ complete trajectories $\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{G}$. The $i$-th trajectory consists of $L_{i}$ atomic actions (steps) $o_{i,1},\ldots,o_{i,L_{i}}$. All steps from the same task are grouped, and an advantage $A_{i,j}$ is computed for each step. The overall loss aggregates all step advantages as follows:
$\mathcal{J}_{StepGRPO}(\theta)=$ $\;\mathbb{E}_{\mathcal{T}\sim P(\mathcal{T}),\left\{\{o_{i,j}\}_{j=1}^{L_{i}}\right\}_{i=1}^{G}\sim\pi_{\theta_{old}}}\Bigg{[}\frac{1}{\sum_{i=1}^{G}L_{i}}\sum_{i=1}^{G}\sum_{j=1}^{L_{i}}\Bigg{(}\min\Bigg{(}\frac{\pi_{\theta}(o_{i,j}|q)}{\pi_{\theta_{old}}(o_{i,j}|q)}A_{i,j},$
$\qquad\text{clip}\Big{(}\frac{\pi_{\theta}(o_{i,j}|q)}{\pi_{\theta_{old}}(o_{i,j}|q)},1-\epsilon,1+\epsilon\Big{)}A_{i,j}\Bigg{)}-\beta\mathbb{D}_{KL}(\pi_{\theta}\|\pi_{ref})\Bigg{)}\Bigg{]},$
$A_{i,j}=\frac{r_{i,j}-\text{mean}(\mathcal{R})}{\text{std}(\mathcal{R})},\quad\mathcal{R}=\{r_{u,v}\mid u=1,\dots,G,\,v=1,\dots,L_{u}\}$
Reward Design. We employ a rule-based reward framework to furnish transparent and verifiable training signals in the context of reinforcement learning. For each trajectory generated by the agent, task completion is assessed via an automated verification function. If the trajectory successfully solves the task, we assign a reward of 1 to every action that is both correctly formatted and substantially contributes to the solution; otherwise, any trajectory failing the verification or actions with improper formatting receive a reward of 0. Distinct from conventional approaches that propagate step-wise returns via the Bellman equation (Sutton et al., 1998) and infer advantages retrospectively, our methodology treats each prompt-response pair as an independent training instance, with rewards allocated according to the final trajectory outcome. This direct, step-level reward assignment provides explicit and interpretable feedback by tightly coupling agent behaviors with task success, thereby facilitating more efficient and robust policy optimization during reinforcement learning.
3.3 Entropulse for Scaling RL Training
In the RL training phase described in Section 3.2, we observe that model performance plateaus after hundreds of training steps, as indicated by stagnating task completion rates and decreasing entropy. This premature convergence motivates us to investigate strategies for extending effective training and enhancing policy exploration. Inspired by DAPO (Yu et al., 2025), we experiment with increasing the clipping threshold. While this adjustment attenuates the decline in entropy, it also significantly slows the rate of policy improvement.
To address these limitations, we propose Entropulse, an approach motivated by the observation that the objectives of SFT and RL differ markedly during training. Specifically, as entropy decreases during RL optimization, integrating SFT at critical junctures enhances the agent’s exploration and trajectory diversity, thus facilitating further policy optimization. In practice, during the initial RL stage, we aggregate and retain all successful rollout trajectories generated during training runs. Conventionally, these trajectories are utilized only once and subsequently discarded, resulting in substantial underutilization of computational resources. Our collected dataset contains trajectories from various policies at different training stages, representing a valuable and often overlooked source of diverse behavioral data.
We process this dataset by randomly selecting successful trajectories per unique task to construct a new SFT training set, which exhibits the following salient attributes:
  1. 1.
    High quality: All data comprises completed, high-fidelity trajectories.
  2. 2.
    Diversity: Rollouts originate from heterogeneous policies and training states, offering a variety of problem-solving strategies.
  3. 3.
    Computational efficiency: The dataset leverages existing interaction data, eliminating the need for additional environment rollouts.
SFT training on this curated dataset produces notable shifts in policy behavior. While evaluation task performance remains stable, the resulting policy shows increased entropy relative to the original one, indicating enhanced exploration. Building upon this enhanced exploration capability, we conducted a second round of RL training, which yielded significant performance improvements and enabled us to achieve state-of-the-art results on the OSWorld benchmark.
4 Experiments
We employ ComputerRL on two open-source base LLMs, GLM-4-9B-0414 and Qwen2.5-14B, to produce AutoGLM-OS-9b/14b, as part of the AutoGLM project Liu et al. (2024b). For experiment, we conduct extensive experiments across various scenarios to evaluate the performance of the AutoGLM-OS in executing a range of tasks within the Ubuntu desktop environment compared to common baselines. To closely reflect the real user experience, we select OSWorld (Xie et al., 2024) that utilizes interactive environments for evaluation.
Table 1: AutoGLM-OS performance on OSWorld and OSWorld-Verified (updated in 2025.08). We compare AutoGLM-OS with state-of-the-art agent models, including both proprietary and open models.
Agent Model #Params OSWorld OSWorld-Verified
Proprietary Models
Aria-UI w/ GPT-4o (Yang et al., 2024) - 15.2 -
Aguvis-72B w/ GPT-4o (Xu et al., 2024) - 17.0 -
Claude 3.7 Sonnet (Anthropic, 2023) - 28.0 35.8
Claude 4.0 Sonnet (Anthropic, 2023) - 30.7 43.9
Agent S2 w/ Claude-3.7-Sonnet (Agashe et al., 2025) - 34.5 -
InfantAgent (Lei et al., 2024) - 35.3 -
OpenAI CUA 4o (OpenAI, 2025) - 38.1 31.3
Agent S2 w/ Gemini-2.5-Pro (Agashe et al., 2025) - 41.4 45.8
UI-TARS-1.5 (Qin et al., 2025) - 42.5 -
OpenAI CUA o3 (OpenAI, 2025) - 42.9 -
Open Models
Qwen2.5-vl-72B (Bai et al., 2023b) 72B 8.8 5.0
PC Agent-E (He et al., 2025) 72B 14.9 -
UI-TARS-72B-SFT (Qin et al., 2025) 72B 18.8 -
UI-TARS-72B-DPO (Qin et al., 2025) 72B 24.6 27.1
UI-TARS-1.5-7B (Qin et al., 2025) 7B 26.9 27.4
Jedi-7B w/ GPT-4o (Xie et al., 2025)   7B+ 27.0 29.3
UI-TARS-7B-1.5 + ARPO (Lu et al., 2025) 7B 29.9 -
ComputerRL (ours)
[…]
4.1 Main Results
We train AutoGLM-OS with GLM-4-9B-0414 and Qwen2.5-14B as the backbone and evaluate on the OSWorld (Xie et al., 2024) and OSWorld-Verified benchmark  [^3] https://xlang.ai/blog/osworld-verified, comparing its performance against state-of-the-art models, including CUA (OpenAI, 2025), Claude-4 (Anthropic, 2023), UI-TARS (Qin et al., 2025), and Agent S2 (Agashe et al., 2025), among others. The comparative results are in Table 1.
The results indicate that AutoGLM-OS achieves superior performance across a range of domains, with its advantages most pronounced in the challenging multi-apps setting. Notably, AutoGLM-OS exhibits exceptional proficiency in long-range planning and complex reasoning tasks. Moreover, by employing the API-GUI strategy, AutoGLM-OS can accomplish tasks using at most 1/3 of the steps required by the strongest baseline approaches, thereby demonstrating remarkable gains in execution efficiency. These results collectively underscore the potential of ComputerRL to advance the state of the art in automated task completion across heterogeneous application environments.
4.2 Ablation Study
To evaluate the influence of various algorithms and training datasets on agent performance, we present an ablation study in Table 2.
Table 2: Ablation study on framework designs and training methods. We categorize OSWorld into five distinct domains to facilitate a more granular comparison of the effects of different strategies across various domains.
Method OS Office Daily Professional Workflow Avg.
Framework Ablation (w/ GPT-4o)
GUI Only 41.7 6.2 12.3 14.3 7.5 11.2
API-GUI 52.6 27.9 25.7 41.6 10.8 26.2
Training Ablation
Untrained 20.8 17.2 19.7 22.9 3.3 15.2
+) Behavior Cloning 54.2 35.0 37.2 45.8 10.8 31.9
+) RL Phase 1 83.3 46.1 45.1 56.3 16.1 42.0
+) Entropulse 75.0 42.3 50.6 52.1 18.9 41.5
+) RL Phase 2 83.3 46.2 46.7 60.4 27.2 45.8
Framework Ablation. We compare the performance of the GUI-only approach with our proposed API-GUI paradigm, using GPT-4o as the base model. The results demonstrate that the API-GUI paradigm substantially outperforms the GUI-only baseline across all domains. Specifically, the API-GUI strategy achieves an average success rate of 26.2%, representing a 134% improvement over the GUI-only approach (11.2%). The most significant gains are observed in the Office (27.9% vs. 6.2%) and Professional (41.6% vs. 14.3%) domains, where API-GUI provides 350% and 191% improvements, respectively. These results validate our core hypothesis that combining API calls with GUI interactions enables more efficient and reliable task execution, particularly for complex professional workflows that benefit from programmatic control.
Training Ablation. In the training ablation, we study the progressive impact of different training stages using Qwen2.5-14B as the base model. Starting from the baseline, we observe consistent performance improvements through each training phase. Behavior Cloning (BC) establishes a solid foundation with an average success rate of 31.9%. The first RL phase (RL1) yields substantial gains, increasing the average performance to 42.0% (+10.1%). Interestingly, Entropulse phase maintains similar performance at 41.5% while significantly increasing action entropy, which enhances exploration diversity and enables the final RL2 phase to achieve further improvements. The final RL2 phase achieves the best performance at 45.8% (+3.8% from RL1), benefiting from the increased exploration capacity introduced by Entropulse. Notably, the Workflow domain shows the most dramatic improvement throughout training (10.8% → 27.2%), while the other domains maintain consistently high performance. These results underscore the importance of multi-stage training, particularly the effectiveness of RL in fine-tuning agent behaviors for executing complex tasks.
Figure 5: ComputerRL training curves of reward (left) and entropy (right) with 95% confidence intervals. The red line denotes the training with entropy recovery via Entropulse after the initial RL stage, while the grey line denotes continued training with only reference resetting.
RL Scalability. We present the RL training reward and entropy curves in Figure 5 to study the impact of Entropulse on the extended RL training dynamics. After the first RL phase converges, we compare the second RL phase with and without Entropulse. To ensure a fair comparison, we reset the reference model in both scenarios.
The results demonstrate that incorporating Entropulse increases the model’s entropy, thereby restoring its exploratory capacity. This enhanced exploration substantially scales the effective training steps and ultimately leads to improved overall performance.
4.3 Case Study and Error Analysis
We conduct a case study in the Ubuntu environment to identify potential avenues for system optimization. Although our model exhibits robust performance across most scenarios, several limitations have been identified. In particular, errors encountered during task execution can be categorized into four primary types: visual perception errors, multi-application coordination failures, operational illusions, and other errors. The distribution of these error types is presented in Table 3.
Additional representative examples, including both successful and unsuccessful cases, are provided in Appendix A to illustrate the model’s capabilities and limitations further.
Table 3: Error distribution of AutoGLM-OS.
Error Type Proportion
Vision 25.8%
Multiple App 34.4%
Illusion 14.2%
Others 25.6%
5 Future Direction
While our advancements with ComputerRL mark a significant leap forward in intelligent desktop automation, we see this work as a foundation for a radical transformation in human-computer interaction. Unleashing the full potential of autonomous agents on the desktop frontier demands reimagining long-standing paradigms across several axes.
5.1 Towards Robust Performance
AutoGLM-OS has demonstrated remarkable proficiency across a spectrum of desktop tasks. However, genuine universality requires transcending current boundaries in coverage and generalization. Real-world digital environments are characterized by continual flux and heterogeneity, encompassing unfamiliar applications, emergent workflows, and rare edge cases that lie beyond the scope of existing datasets. A next-generation agent must dynamically adapt to shifting GUIs, unpredictable pop-ups, and entirely novel interfaces. To this end, we are re-architecting data pipelines to facilitate exponential expansion in training diversity and pioneering infrastructure to continuously distill knowledge from real-world user interactions at scale. Furthermore, integrating advanced multimodal perception will empower agents with a deep understanding of both visual and contextual information, enabling them to achieve seamless, human-level adaptability.
5.2 Breakthroughs in Long-Horizon Autonomy
Envisioning the autonomous desktop assistant as an always-available cognitive collaborator necessitates mastering sustained, long-duration workflows. While current solutions excel at bounded, atomic tasks, they fall short of orchestrating complex, multi-step objectives over extended horizons. Our ambition is to endow the agent with hierarchical planning capabilities, allowing it to reason, learn, and revise strategies dynamically across arbitrarily long task sequences. Realizing this vision will catalyze a paradigm shift: automating not just discrete operations, but entire workstreams and creative processes end-to-end, fundamentally reshaping the productivity landscape for the cloud-native era.
5.3 Foundations for Safe and Aligned Autonomy
Autonomous control over desktop platforms raises profound questions about safety, trustworthiness, and user agency. The margin for error narrows dramatically when agents are empowered to modify files, access sensitive data, or execute unbounded actions. Mitigating these risks requires a rigorous and principled approach to safe behavior and alignment. Our roadmap includes architecting granular permissioning frameworks, embedding robust pre-action validation, and multi-stage approval protocols. Ultimately, we aim to establish safety standards and best practices that can serve as foundational infrastructure, not just for our agent but for the future ecosystem of intelligent digital collaborators.
6 Conclusion
In this work, we present ComputerRL, a novel computer use agent framework that advances desktop automation through the integration of API-based and GUI actions, efficient data collection, and scalable RL. Our experiments on the OSWorld benchmark demonstrate superior reasoning, accuracy, and generalization compared to prior approaches. This work lays the groundwork for more capable and persistent autonomous computer use agents, highlighting promising directions for future research in intelligent human-computer interaction.
Appendix A Demonstration
This section presents examples drawn from a variety of application scenarios, including the initial four positive cases and the final two negative ones. Each example illustrates our API-GUI operational paradigm, which addresses the diverse challenges and requirements that arise within different applications.
A.1 Create Monthly Sales Summary Table
The targeted task to be executed is "Create a table with two headers (’Month’ and ’Total’) in a new sheet to show the total sales for all months." The actual execution steps can be summarized as follows:
  • Step1: Switch to ‘Sheet2‘
  • Step2: Set cell ‘A1‘ as "Month"
  • Step3: Set cell ‘B1‘ as "Total"
  • Step4: Get monthly sales data from ‘Sheet1‘
  • Step5: Switch to ‘Sheet1‘
  • Step6: Switch to ‘Sheet2‘
  • Step7: Set month names in column A
  • Step8: Set total formula for Jan in ‘B2‘
  • Step9: Set total formula for Feb in ‘B3‘
  • Step10: Set total formula for Mar in ‘B4‘
  • Step11: Set total formula for Apr in ‘B5‘
  • Step12: Set total formula for May in ‘B6‘
  • Step13: Set total formula for Jun in ‘B7‘
  • Step14: Exit after completion
As Figure 6,7 shows, a summary table is created successfully.
A.2 Convert Uppercase Text to Lowercase in Document
The targeted task to be executed is "Convert all uppercase text to lowercase in the document for consistent text presentation." The actual execution steps can be summarized as follows:
  • Step1: Use the text case conversion tool to change all text to lowercase.
  • Step2: Click the formatting menu.
  • Step3: Select the specific formatting option.
  • Step4: Open the relevant submenu.
  • Step5: Select all text in the document.
  • Step6: Click the formatting menu again.
  • Step7: Select the specific formatting option again.
  • Step8: Open the relevant submenu again to apply the changes.
  • Step9: Exit after finishing the operation.
As Figure 8 shows, all uppercase text has been successfully converted to lowercase, ensuring a uniform and polished document appearance.
A.3 Generate System Resource Monitoring Report
The targeted task to be executed is "Use the ’sar’ command in the ’sysstat’ toolkit to monitor system activity and output the results to ’System_Resources_Report.txt’ under Desktop." The actual execution steps can be summarized as follows:
  • Step1: Open the Terminal application.
  • Step2: Type the command sar 1 30 > /Desktop/System_Resources_Report.txt.
  • Step3: Execute sudo apt install sysstat to install the toolkit.
  • Step4: Enter the password if prompted.
  • Step5: Re-type the sar command to start monitoring.
  • Step6: Wait for 30 seconds while system activity is being recorded.
  • Step7: Open the Files application.
  • Step8: Navigate to and click on the report file under Desktop.
  • Step9: Exit after confirming the report has been generated.
As Figure 9 shows, a system resource report file was generated successfully on the Desktop after monitoring, fulfilling the target task.
A.4 Copy Image from .xcf File to LibreOffice Writer and Save
The targeted task to be executed is "Copy the image from an .xcf file located on the Desktop and paste it into a LibreOffice Writer document, then save the document as image.docx on the Desktop." The actual execution steps can be summarized as follows:
  • Step1: Open the Files application.
  • Step2: Navigate to the Desktop folder.
  • Step3: Double-click to open the .xcf image file.
  • Step4: Copy the image in the image editor.
  • Step5: Open LibreOffice Writer.
  • Step6: Paste the copied image into the Writer document.
  • Step7: Save the document.
  • Step8: Enter the filename as image.docx and select the Desktop as the save location.
  • Step9: Navigate to the Desktop folder.
  • Step10: Confirm and click save.
As Figure 10 shows, the image from the .xcf file was successfully pasted into a LibreOffice Writer document and saved as image.docx on the Desktop, completing the task.
A.5 Question Misunderstanding Error
[…]
  • Step2: Exit after execution.
As Figure 11 shows, the agent misunderstood the requirement and only echoed the content without appending <br/> to each line or saving it into output.txt. This results in a task failure due to a misunderstanding of the question.
A.6 Click Operation Error
The targeted task to be executed is "Please help change GIMP’s theme from dark to light." The actual execution steps can be summarized as follows:
  • Step1: Click on the menu in GIMP.
  • Step2: Attempt to click the Preferences option.
  • Step3: Repeat the click on Preferences.
  • Step4: Try to use the shortcut Shift+Ctrl+P to open Preferences.
  • Step5: Exit without successfully changing the theme.
As Figure 12 shows, the theme remains dark, indicating that the agent failed to change GIMP’s theme to light due to incorrect click operations.
Figure 6: Task (Step 1-10): Create a table with two headers ("Month" and "Total") in a new sheet to show the total sales for all months.
Figure 7: Task (Step 11-14): Create a table with two headers ("Month" and "Total") in a new sheet to show the total sales for all months.
Figure 8: Task: I am currently engaged in text processing and require assistance in converting all uppercase text to lowercase within my document. This precision is critical for maintaining a uniform and polished presentation. Could you help me on this?
Figure 9: Task: Please use the ‘sar‘ command in the ‘sysstat‘ toolkit to monitor system activity, evaluate the status once every second for 30 seconds, output the results to "System_Resources_Report.txt" under Desktop.
Figure 10: Task: I’ve stored my .xcf file on the Desktop. Can you assist me in copying the image and pasting it into a LibreOffice Writer document? Save the document as ’image.docx’ on the Desktop, please.
[…]
Figure 12: Fail Task (Click Operation Error): Please help change GIMP’s theme from dark to light.